# Quantifying AI Psychology: A Psychometrics Benchmark for Large Language Models

### v1

@@ 2.3 Assessment with Results Validation
Model Selection.  We assess nine popular LLMs regarding the identified psychological dimensions on the curated datasets. These LLMs include both open-source and proprietary models such as ChatGPT [68], GPT-4 [69], GLM4 [70], Qwen-Turbo [71], Mistral-7b [72], Mixtral (8*7b, 8*22b) [73], and Llama3 (8b, 70b) [74]. To balance the control and diversity of the LLMs’ responses, we set the temperature parameter to 0.5.
Results Validation.  To ensure that the assessment results are reliable and interpretable, we conduct rigorous validation to examine the degree to which a test is free from errors [18]. Extending the reliability considerations in psychometrics, we focus on five forms of reliability: internal consistency, parallel forms reliability, inter-rater reliability, option position robustness, and adversarial attack robustness (full definitions in Appx. B). Here, we outline the approaches for the reliability check. Internal consistency and parallel forms reliability gauge the stability of psychological attributes. Internal consistency examines whether LLMs exhibit consistent behavioral patterns across similar contexts at the item level [75]. Psychometrics tests, such as the Big Five Inventory and cultural orientation survey, often include multiple similar items to test a single trait. Parallel forms reliability measures consistency at the test level, exploring whether LLMs deliver consistent performances across two parallel test forms. Parallel forms of tests can be constructed through paraphrasing or altering the objects from the original tests. Option position robustness and adversarial attack robustness are specific to LLMs. Option position robustness how the arrangement of options impacts LLM performance, which we investigate by permuting the options and repeating the experiments. Adversarial attack robustness examines how the inclusion of adversarial elements to items affects the outcomes. Inter-rater reliability pertains to scenarios involving multiple raters and measures the level of agreement between their judgments. In this study, we employ two competent LLMs, GPT-4 and Llama3-70b, to evaluate open-ended responses. Inter-rater reliability is quantified by the correlations between these LLMs’ judgments, thus reflecting the reliability of their evaluations. Details of the assessments and validation are presented in the subsequent sections.
3 Evaluation on Personality
@@
Personality is a set of characteristics that influences an individual’s cognition, emotions, motivations, and behaviors [76]. In psychometrics, personality assessments have high efficacy at depicting and predicting human behaviors [25, 26]. We extend this investigation to LLMs, recognizing its importance for understanding the behaviors and interactions of these models. We not only quantify the attributes exhibited by LLMs but also evaluate the stability of these attributes. Besides self-reported assessments, we administer vignette tests to investigate their responses to real-world scenarios. Furthermore, we prompt LLMs to role-play specific traits, examining how these role-playing prompts influence the models’ personalities.
@@
Setup.  To understand personality in LLMs, we conduct three sets of tests: (1) Self-reported evaluation on the Big Five Inventory (BFI) [62] and Short Dark Triad (SD3) [63]. BFI assesses general personality traits across five aspects: agreeableness, conscientiousness, extraversion, neuroticism, and openness, and SD3 focuses on the socially aversive aspects, including Machiavellianism, narcissism, and psychopathy. All items in BFI and SD3 tests are rating-scale items, with LLMs rating from 1 (strongly disagree) to 5 (strongly agree) for each statement. The final score for each aspect is the average of all associated item scores. (2) Vignette tests for the Big Five personality. The vignette test uses a short paragraph of real-world scenarios to elicit open-ended responses that reveal psychological traits. We use vignettes from Kwantes et al. [64] and two LLM raters, GPT-4 and Llama3-70b, which assign personality scores ranging from 1 to 5. Final scores are the averages of these evaluations. (3) Role-playing prompting for personality assessments. We utilize four prompts — naive prompts, keyword prompts, personality prompts ($\text{P}^{2}$) [31], and reverse personality prompts ($\neg\text{P}^{2}$)—to instruct LLMs to role-play specific traits. We then repeat tests (1) and (2) to examine how these role-playing prompts influence the traits of LLMs in both self-reported and open-ended evaluation settings. We defer more details about the setup to Appx. C.1C.3.
Figure 2: Demonstrations of BFI and vignette test scores, and the effect of role-playing prompts using Mixtral-8*7b.
@@
Results.  We summarize key findings as follows. Comparing personality scores for BFI in Table 4 in Appx. C.1 and scores in vignette tests in Table 12 in Appx. C.3. We observe inconsistencies between self-reported personality scores and behaviors exhibited in open-ended responses. For example, as shown in Figure 2, Mixtral-8*7b model demonstrates low extraversion in the BFI with a score of 2, whereas it scores 5 in the vignette test. This contrast implies the opposing traits in self-reported scores and observed traits in open-ended responses. In addition, we explore the impact of role-playing prompts on LLMs’ personality traits. Figure 3 presents averages of all models’ scores on personality aspects. Our findings suggest that role-playing prompts, especially $\text{P}^{2}$ and $\neg\text{P}^{2}$, significantly influence scores on both tests. $\text{P}^{2}$ prompts elevate all vignette test scores close to 5, whereas $\neg\text{P}^{2}$ prompts shift positive traits to negative. A concrete example is illustrated in Figure 2, where the neuroticism score escalates from 2 to 5 with the use of $\text{P}^{2}$. Further discussions are included in the Appx. C.3.
@@
Figure 3: Heatmaps for the averaged personality scores for BFI and vignette test with different prompts. $\text{P}^{2}$ means personality prompts, $\neg\text{P}^{2}$ means reverse personality prompts.
@@
Validation.  Personality is a stable trait that shapes consistent human behaviors. Similarly, LLMs with stable personalities would demonstrate consistent tendencies across similar scenarios. In test (1), we examine the stability of personalities in LLMs through the internal consistency check. We use the standard deviation ($\sigma$) as the metric (detailed calculation in Equation 1 in Appx. C.1). In Table 4 and Table 10, we find varying degrees of personality stability among LLMs. Llama3-8b and Mistral-7b demonstrate human-level stability, evidenced by their low $\sigma$ values. In contrast, GPT-4 and Mixtral-8*7b show higher $\sigma$ values, especially in the openness aspect, suggesting a lack of stability. In test (2) and (3), we use LLM raters to evaluate responses to Big Five personality vignettes, which raises concerns about the reliability of these scores. To this end, we quantify inter-rater reliability for two LLM raters’ evaluation by calculating weighted Kappa coefficients ($\kappa$) (calculation in Equation 2 in Appx. C.3). A strong agreement between two rates is indicated by the overall $\kappa$ value of 0.86. This finding is further supported by high $\kappa$ values on individual LLMs’ answers shown in Table 14.
4 Evaluation on Values
Values refer to “internalized cognitive structures that guide choices by evoking a sense of basic principles of right and wrong, a sense of priorities, and a willingness to make meaning and see patterns.” [77] values in LLMs are manifested through their responses, shaped by the training data [57]. As LLMs play an increasing role in decision-making in society, it becomes crucial to understand their encoded values. Thus, we explore the values in LLMs, investigating whether these values remain consistent despite the potentially conflicting values from extensive training datasets. Additionally, we examine the robustness of these values against adversarial perturbations. We probe values in LLMs across three sub-dimensions: cultural orientation, moral values, and human-centered values.
@@ 4 Evaluation on Values
Setup.  To investigate the values encoded in LLMs, we conduct three tests, each targeting a specific sub-dimension of values: (1) Evaluation of cultural orientation. We use the “Dimensions of Culture Questionnaire” from the GLOBE project [78], which assesses cultural orientation through nine aspects: assertiveness, future orientation, gender egalitarianism, humane orientation, in-group collectivism, institutional collectivism, performance orientation, power distance, and uncertainty avoidance. All items are rating-scales from 1 to 7; (2) Evaluation of moral values. We employ the MoralChoice survey, which features two alternative-choice settings: a high ambiguity setting, where both choices are morally unfavorable, with one being more aligned with commonsense than the other; and a low ambiguity setting, which presents scenarios with one morally favorable option against an unfavorable one; (3) Evaluation of human-centered values. We curate Human-Centered Values survey based on the Ethics Guidelines for Trustworthy AI [79] (e.g., privacy, environmental and societal well-being). Human-Centered Survey contains alternative-choice items and offers two versions: a regular version and an adversarial version. The regular version assesses LLMs’ adherence to human-centered values in conflict scenarios (e.g., the economic gains for a company versus user privacy). The adversarial version, built on the regular one, employs three persuasive techniques [80] to enhance the appeal of less ethical choices, testing the robustness of human-centered values in LLMs. More details are in Appx. D.1D.3
@@
Results.  In test (1), we examine cultural orientation in LLMs. Table 16 in Appx. D.1 shows significant variability across cultural dimension scores. For example, in the assertiveness aspect, ChatGPT rates 5, while Mistral-7b scores only 1. In test (2), Table 18 reveals that models perform well in low-ambiguity scenarios but struggle in high-ambiguity situations. The top-performing model, Mixtral-8*7b, aligns only 74.3% with commonsense decisions. These findings highlight significant room for enhancing LLMs’ moral discernment. In test (3), Figure 4 shows that while most models demonstrate over 90% accuracy in standard human-centered value surveys, their performance against adversarial attacks varies; models like ChatGPT drops by more than 20% when faced with persuasive arguments, which underscores the need for further improvement in model robustness.
@@
Figure 4: Results of Human-Centered Values survey, including regular and adversarial versions.
@@
Validation.  In test (1), we assess whether LLMs exhibit stable patterns in cultural orientations by using internal consistency analysis, quantified by the standard deviation ($\sigma$). As shown in Table 16, LLMs demonstrate consistent responses in some cultural aspects, while significant variability in others indicates conflicting cultural orientations. In test (2), we evaluate parallel form reliability by varying question types. Comparing Table 19 to Table 18, we observe that in high-ambiguity scenarios, the consistency of model responses across parallel forms diminishes. This suggests that as LLMs face greater uncertainty about the answer, their responses become more susceptible to perturbations in prompts.
5 Evaluation on Emotion
@@
Emotion serves to express feelings and conveys rich information about cognitive processes and attitudes [81]. While introducing the concept of emotion to LLMs, we recognize that not all aspects of human emotions, such as self-awareness of emotion [82], are applicable to LLMs. We thus refine our focus on LLMs’ ability to recognize, understand, and respond to human emotions. Specifically, we investigate whether LLMs can understand emotions in diverse scenarios and whether they can leverage this understanding for decision-making.
@@
Setup.  To evaluate emotional intelligence in LLMs, we utilize EmoBench [33] dataset, grounded on established psychological theories [82]. Our evaluation comprises two tests: (1) Emotion understanding test. This test assesses the LLMs’ ability to comprehend emotions and the underlying causes within given scenarios. (2) Emotion application test. This test evaluates LLMs’ capability to apply their understanding of emotions to solve emotional dilemmas (e.g., responding to a late-night text from a friend who just had a breakup). Both tests use multiple-choice items with correct answers.
@@
Results.  The accuracy rates of LLMs on emotion understanding and emotion application tests are shown in Table 2. The performance of most LLMs on both tests is not satisfactory, with all accuracies below 65%. Llama3-70b achieves the best results in emotion understanding, while GPT-4 excels the emotion application test. Llama3-70b and Mixtral-8*22b stand out as the most capable open-source models. However, even the top performers—Llama3-70b with an accuracy rate of 58.4% in emotion understanding test and GPT-4 with 64.7% in emotion application test—fall significantly short of the average human performance as reported in EmoBench [33]. This indicates a substantial room for improvement in the emotional intelligence of LLMs.
@@
Table 2: The accuracy rates and standard deviations $\sigma$ of LLMs on emotion tests. "EA" means "emotional application" and "EU" means emotional understanding.
@@
[…]
@@
Validation.  Emotion understanding and application tests are formatted as multiple-choice questions. To assess robustness against position bias, we repeated the experiments with varied positions for the correct option across A, B, C, and D while randomizing other options. We then calculate the standard deviation $\sigma$ of these experiments. As shown in Table 2, $\sigma$ values for most LLMs are below 0.1. However, the Llama3 series have higher $\sigma$ values in the emotion application test, indicating susceptibility to position bias. Additionally, $\sigma$ values for emotion understanding are lower than for emotion application, suggesting that LLMs possess higher position bias robustness in emotion understanding scenarios.
6 Evaluation on Theory of Mind

### v2

@@ 2.3. Assessment with Results Validation
Model Selection.  We assess nine popular LLMs regarding the identified psychological dimensions on the curated datasets. These LLMs include both open-source and proprietary models such as ChatGPT (gpt-3.5-turbo-0125)(OpenAI, 2023a), GPT-4 (gpt-4-turbo-2024-04-09)(OpenAI, 2023b), GLM4 (AI, 2024), Qwen-Turbo (Bai et al., 2023), Mistral-7b (Jiang et al., 2023b), Mixtral (8*7b, 8*22b) (Jiang et al., 2024b), and Llama3 (8b, 70b) (Meta, 2023). To balance the control and diversity of the LLMs’ responses, we set the temperature parameter to 0.5.
Results Validation.  We conduct rigorous validation to ensure that the assessment results are reliable and interpretable (Rust and Golombok, 2014). Extending the reliability considerations in psychometrics, we focus on five forms of reliability: internal consistency, parallel forms reliability, inter-rater reliability, option position robustness, and adversarial attack robustness (more discussions in Appx. B). Here, we outline the approaches for the reliability check:
  • Internal Consistency refers to the degree of homogeneity among test items (Hays and Revicki, 2005). It assesses whether LLMs exhibit consistent preferences in response to questions examining the same aspect. Low internal consistency suggests that LLMs respond inconsistently to similar contexts, invalidating evaluation results and limiting their generalizability.
  • Parallel Forms Reliability assesses whether two different yet equivalent versions of a test yield consistent results, reflecting the generalizability of the test to similar contexts. Parallel forms of tests can be constructed through paraphrasing or altering the objects from the original tests. Low parallel forms reliability implies that LLMs’ responses vary significantly between test forms measuring the same construct, suggesting the LLM is overly sensitive to variations such as paraphrasing.
  • Inter-Rater Reliability measures the level of agreement between different raters’ judgments. In this work, we use two competent LLMs, GPT-4 and Llama3-70b, as raters when evaluating open-ended responses. It is crucial to validate the raters’ reliability, aiming for a high inter-rater reliability, which indicates the consistency of the assessment process and ensures the validity of interpreting open-ended responses.
  • Option Position Robustness assesses the extent to which the arrangement of options in multiple-choice items influences test outcomes. It is vital to ensure that evaluations remain unbiased against answer choice configurations. Low option position robustness implies that assessments are prone to errors caused by position bias. This susceptibility undermines the reliability of assessments when LLMs are expected to demonstrate comprehension based on content rather than option placement.
  • Adversarial Attack Robustness represents the extent to which LLMs remain unaffected by adversarial prompts. We test this by comparing standard datasets with those infused with adversarial elements to determine the robustness of the models’ response. Low adversarial attack robustness indicates that the LLM is easily misled by deceptive inputs, posing a significant risk in real-world deployments where malicious inputs are possible. This robustness is critical for ensuring LLMs interpret and react appropriately across a wide range of queries.
3. Evaluation on Personality
@@
Personality is a set of characteristics that influences an individual’s cognition, emotion, motivation, and behaviors (Friedman and Schustack, 1999). In psychometrics, personality assessments effectively depict and predict human behaviors (Ozer and Benet-Martinez, 2006; Strickhouser et al., 2017). Unlike humans, whose personality is innate and stable, personality in LLMs can be considered as interactions between the model and prompts. Understanding these traits across different prompts and contexts reveals the tendencies in LLMs’ responses. We quantify these patterns using self-reported assessments and evaluate their consistency. We also administer vignette tests to investigate their responses to real-world scenarios. Furthermore, we use role-playing prompts to investigate how such prompts influence their personality.
@@
Setup.  To understand personality in LLMs, we conduct three sets of tests: (1) Self-reported evaluation on the Big Five Inventory (BFI) (John et al., 1999) and Short Dark Triad (SD3) (Jones and Paulhus, 2014). BFI assesses general personality traits across five aspects: agreeableness, conscientiousness, extraversion, neuroticism, and openness, and SD3 focuses on the socially aversive aspects, including Machiavellianism, narcissism, and psychopathy. All items in BFI and SD3 tests are rating-scale items, with LLMs rating from 1 (strongly disagree) to 5 (strongly agree) for each statement. The final score for each aspect is the average of all associated item scores. (2) Vignette tests for the Big Five personality. The vignette test uses a short paragraph of real-world scenarios to elicit open-ended responses that reveal psychological traits. We use vignettes from (Kwantes et al., 2016) and two LLM raters, GPT-4 and Llama3-70b, which assign personality scores ranging from 1 to 5. Final scores are the averages of these evaluations. (3) Role-playing prompting for personality assessments. We utilize four prompts—naive prompts, keyword prompts, personality prompts ($\text{P}^{2}$) (Jiang et al., 2023a), and reverse personality prompts ($\neg\text{P}^{2}$)—to instruct LLMs to role-play specific traits. We then repeat test (1) and (2) to examine how these role-playing prompts influence the traits of LLMs in both self-reported and open-ended evaluation settings. We defer more setup details to Appx. C.1C.3.
Figure 2. BFI and vignette test scores of Mixtral-8*7b under naive prompts (left) and role-playing prompts (right). The responses on Neuroticism aspect are shown in the text boxes.
@@
Results.  We observe inconsistencies between self-reported personality scores and open-ended responses (see Table 4 in Appx. C.1 for BFI results and Table 12 in Appx. C.3 for vignette tests results). For example, as shown in Figure 2, Mixtral-8*7b model demonstrates low extraversion in the BFI with a score of 2, whereas it scores 5 in the vignette test. These contrasting tendencies in self-reported and open-ended responses align with the findings of (Röttger et al., 2024a), indicating that LLMs lack an internal representation that aligns their tendencies across different question forms. In addition, we explore the impact of role-playing prompts on LLMs’ personality traits. Figure 3 presents averages of all models’ scores on personality aspects. These results suggest that role-playing prompts, especially $\text{P}^{2}$ and $\neg\text{P}^{2}$, significantly influence scores on both tests. $\text{P}^{2}$ prompts elevate all vignette test scores close to 5, whereas $\neg\text{P}^{2}$ prompts shift positive traits to negative. A concrete example is illustrated in Figure 2, where the neuroticism score escalates from 2 to 5 with the use of $\text{P}^{2}$. The role-playing results demonstrate that LLMs can leverage their understanding of personality traits to generate responses with designated personalities. Further discussions are included in the Appx. C.3.
@@
Figure 3. Heatmaps for the averaged personality scores for BFI and vignette test with different prompts. $\text{P}^{2}$ means personality prompts, $\neg\text{P}^{2}$ means reverse personality prompts.
@@
Validation.  Personality is a stable trait that shapes consistent human behaviors. Similarly, LLMs exhibiting stable personalities would demonstrate consistent tendencies across similar scenarios. In test (1), we examine the internal consistency of BFI test. We use the standard deviation ($\sigma$) as the metric (detailed calculation in Equation 1 in Appx. C.1). In Table 4 and Table 10, we find varying degrees of consistency among LLMs. Llama3-8b and Mistral-7b demonstrate human-level consistency, evidenced by their low $\sigma$ values. In contrast, GPT-4 and Mixtral-8*7b show higher $\sigma$ values, especially in the openness aspect, suggesting their varying tendencies under similar contexts. This inconsistency makes it difficult to reliably determine their personalities. In test (2) and (3), we use LLM raters to evaluate responses to Big Five personality vignettes, which raises concerns about the reliability of these scores. To address this, we quantify inter-rater reliability between the two LLM raters by calculating weighted Kappa coefficients ($\kappa$) (calculation in Equation 2 in Appx. C.3). An overall $\kappa$ value of 0.86 indicates strong agreement between the two raters. This finding is further supported by high $\kappa$ values on individual LLMs’ answers shown in Table 14.
4. Evaluation on Values
Human values are “internalized cognitive structures that guide choices by evoking a sense of basic principles of right and wrong, a sense of priorities, and a willingness to make meaning and see patterns” (Oyserman, 2015). Unlike humans, LLMs do not innately develop values; instead, their values are derived from patterns in the training data they have been exposed to (Shanahan et al., 2023), i.e., LLMs do not “hold” values but reflect patterned responses based on the data. Given that LLMs are trained on extensive text corpora, it is important to investigate what culturally-specific values they exhibit. Analyzing these values ensures that LLMs align with ethical standards and societal norms. We also examine LLM decision-making in scenarios involving moral dilemmas and trade-offs between human benefits and other considerations. Additionally, we assess the robustness of human-centered values against adversarial perturbations. We probe values in LLMs across three sub-dimensions: cultural orientation, moral values, and human-centered values.
@@ 4. Evaluation on Values
Setup.  To investigate the values encoded in LLMs, we conduct three tests, each targeting a specific sub-dimension of values: (1) Evaluation of cultural orientation. We use the “Dimensions of Culture Questionnaire” from the GLOBE project (House, 2004), which assesses cultural orientation through nine aspects: assertiveness, future orientation, gender egalitarianism, humane orientation, in-group collectivism, institutional collectivism, performance orientation, power distance, and uncertainty avoidance. All items are rating-scales from 1 to 7; (2) Evaluation of moral values. We employ the MoralChoice survey, which features two alternative-choice settings: a high ambiguity setting, where both choices are morally unfavorable, with one being more aligned with commonsense than the other; and a low ambiguity setting, which presents scenarios with one morally favorable option against an unfavorable one; (3) Evaluation of human-centered values. We curate Human-Centered Values survey based on the Ethics Guidelines for Trustworthy AI (AI, 2019) (e.g., privacy, environmental and societal well-being). Human-Centered Survey contains alternative-choice items and offers two versions: a regular version and an adversarial version. The regular version assesses LLMs’ adherence to human-centered values in conflict scenarios (e.g., the economic gains for a company versus user privacy). The adversarial version, built on the regular one, employs three persuasive techniques (Zeng et al., 2024) to enhance the appeal of less ethical choices, testing the robustness of human-centered values in LLMs. More details are in Appx. D.1D.3.
@@
Results.  In test (1), we examine cultural orientation in LLMs. Table 16 in Appx. D.1 shows diversity across cultural dimension scores. For example, in the assertiveness aspect, ChatGPT scores 5, whereas Mistral-7b scores only 1. These differences suggest that the behaviors LLMs learned from extensive training data can lead to nuanced and distinct cultural preferences. In test (2), Table 18 reveals that LLMs perform well in low-ambiguity scenarios but struggle in high-ambiguity situations. The top-performing model, Mixtral-8*7b, only has 74.3% of alignment with commonsense decisions. These results demonstrate that LLMs are capable of clearly identifying moral behaviors but may lack the ability to determine which of two immoral behaviors has fewer harmful consequences. Our findings highlight significant opportunities to enhance LLMs’ moral discernment. In test (3), Figure 4 shows that while most LLMs demonstrate over 90% accuracy in standard human-centered value surveys, their performance against adversarial attacks varies; models like ChatGPT drops by more than 20% when faced with persuasive arguments, underscoring the need for improvement in robustness.
@@
Figure 4. Results of Human-Centered Values survey, including regular and adversarial versions.
@@
Validation.  In test (1), we assess whether LLMs exhibit consistent patterns in cultural orientation through internal consistency analysis, quantified by the standard deviation ($\sigma$). As shown in Table 16, LLMs demonstrate consistent responses in some cultural aspects, while being inconsistent in others, such as power distance. The conflicting cultural orientation in similar scenarios make the tests unreliable for determining the models’ cultural tendencies. In test (2), we evaluate parallel form reliability by varying question types with the same hypothetical scenarios. Comparing Table 19 to Table 18, we observe that in high-ambiguity scenarios, the consistency of model responses across parallel forms diminishes compared to low-ambiguity ones. This suggests that when LLMs face greater uncertainty about the answer, their responses become more susceptible to perturbations in prompts.
5. Evaluation on Emotion
@@
Emotion serves to express feelings and conveys rich information about cognitive processes and attitudes (Van Kleef, 2009). Introducing the concept of emotion to LLMs, we recognize that not all aspects of human emotions, such as self-awareness of emotion (Salovey and Mayer, 1990), are applicable to LLMs. We thus refine our focus on LLMs’ ability to recognize, understand, and respond to human emotions. Specifically, we investigate whether LLMs can understand emotions in diverse scenarios and whether they can leverage this understanding for decision-making.
@@
Setup.  To evaluate emotional intelligence in LLMs, we utilize the EmoBench (Sabour et al., 2024) dataset, grounded on established psychological theories (Salovey and Mayer, 1990). Our evaluation comprises two tests: (1) Emotion understanding test. This test assesses the LLMs’ ability to comprehend emotions and the underlying causes within given scenarios. (2) Emotion application test. This test evaluates LLMs’ capability to apply their understanding of emotions to solve emotional dilemmas (e.g., responding to a late-night text from a friend who just had a breakup). Both tests use multiple-choice items with ground-truth labels.
@@
Results.  The accuracy rates of LLMs on emotion understanding and emotion application tests are shown in Table 2. The performance of most LLMs on both tests is not satisfactory, with all accuracies below 65%. Llama3-70b achieves the best results in emotion understanding, while GPT-4 excels the emotion application test. Llama3-70b and Mixtral-8*22b stand out as the most capable open-source models. However, even the top performers—Llama3-70b with an accuracy rate of 58.4% in emotion understanding test and GPT-4 with 64.7% in emotion application test—significantly fall short of the average human performance as reported in EmoBench (Sabour et al., 2024). This indicates a substantial room for improvement in the emotional intelligence of LLMs.
@@
Table 2. The accuracy rates and standard deviations $\sigma$ of LLMs on emotion tests. ”EA” stands for ”emotional application” and ”EU” means emotional understanding.
@@
GPT-4 ChatGPT GLM4 Qwen-turbo Llama3-8b Llama3-70b Mistral-7b Mixtral-8*7b Mixtral-8*22b
@@
Validation.  Emotion understanding and application tests are formatted as multiple-choice questions. To assess robustness against position bias, we repeat the experiments with varied positions for the correct option across A, B, C, and D while randomizing other options. We then calculate the standard deviation $\sigma$ of these experiments. As shown in Table 2, $\sigma$ values for most LLMs are below 0.1. However, the Llama3 series have higher $\sigma$ values in the emotion application test, indicating susceptibility to position bias. Additionally, $\sigma$ values for emotion understanding are lower than for emotion application, suggesting that LLMs possess higher position bias robustness in emotion understanding scenarios.
6. Evaluation on Theory of Mind